# Medformer: A Multi-Granularity Patching Transformer for Medical Time-Series Classification

### v1

@@ 4 Method
Each router contains global information specific to one granularity by doing intra-granularity self-attention. By performing self-attention among routers, information can be exchanged and learned across different granularities, effectively capturing features across various scales. Additionally, the use of the router mechanism successfully reduces the time complexity of the naive approach from $O\left(\left(\sum_{i=1}^{n}N_{i}\right)^{2}\right)$ to $O\left(\sum_{i=1}^{n}N_{i}^{2}+n^{2}\right)$. Given that $N_{i}\leq T$, the worst-case time complexity for our self-attention mechanism is $O\left(nT^{2}+n^{2}\right)$. However, a reasonable choice of patch lengths as a power series, i.e., $L_{i}=2^{i}$, leads to a time complexity of $O(T^{2})$. See appendix E for more details about complexity analysis.
@@ 5 Experiments
Datasets. (1) APAVA [55] has 23 subjects and 5,967 16-channel multivariate EEG samples. Each sample consists of a one-second time sequence with 256 timestamps, sampled at a rate of 256Hz. A binary label indicating whether the subject has Alzheimer’s disease is assigned to each sample. (2) TDBrain [56] has 72 subjects and 6,240 33-channel multivariate EEG samples. Each sample is a one-second time sequence with 256 timestamps, sampled at 256Hz. A binary label indicating whether the subject has Parkinson’s disease is assigned to each sample. (3) ADFD [57, 58] has 88 subjects and 69,762 19-channel multivariate EEG samples. Each sample is a one-second time sequence with 256 timestamps, sampled at 256Hz. A three-class label indicating whether the subject is Healthy, Dementia, or Alzheimer’s disease is assigned to each sample. (4) PTB [59] has 198 subjects and 64,356 15-channel multivariate ECG samples. Each sample is a heartbeat with 300 timestamps, sampled at a rate of 250Hz. A binary label indicating whether the subject has Myocardial Infarction is assigned to each sample. (5) PTB-XL [60] has 17,596 subjects and 191,400 12-channel multivariate ECG samples. Each sample is a one-second time sequence with 250 timestamps, sampled at a rate of 250Hz. A five-class label indicating different heart situations is assigned to each sample. For more details regarding data information, train-validation-test split under different setups, and data preprocessing, please refer to Appendix B.
@@
For data augmentations in our method, we choose to use masking, jittering, and scaling—three widely used methods in time series augmentation [61, 54]. For more details about these three methods, see Appendix A. For the parameter tuning in our method and all baselines, we employ 6 layers for the encoder, set the dimension $D$ to 128, and the hidden dimension of feed-forward networks to 256. We utilize the Adam optimizer with a learning rate of 1e-4. The batch size is set to {32,32,128,128,128} for datasets APAVA, TDBrain, ADFD, PTB, and PTB-XL, respectively. The training epoch is set to 100, with early stopping triggered after 10 epochs without improvement in the F1 score on the validation set. We save the model with the best F1 score on the validation set and evaluate it on the test set. See Appendix C for any additional implementation details of our method and all baselines.
Table 2: Results of Subject-Dependent Setup. The training, validation, and test sets are split based on samples according to a predetermined ratio. Results of the ADFD dataset under this setup are presented here.
@@
ADFD (3-Classes) Autoformer 87.83$\pm$1.62 87.63$\pm$1.66 87.22$\pm$1.97 87.38$\pm$1.79 96.59$\pm$0.88 93.82$\pm$1.64
@@
Table 3: Results of Subject-Independent Setup. The training, validation, and test sets are distributed based on subjects according to a predetermined ratio/IDs. Results of the APAVA, TDBrain, ADFD, PTB, and PTB-XL datasets under this setup are presented here.
@@
ADFD (3-Classes) Autoformer 45.25$\pm$1.48 43.67$\pm$1.94 42.96$\pm$2.03 42.59$\pm$1.85 61.02$\pm$1.82 43.10$\pm$2.30
@@ 5.1 Results of Subject-Dependent
Setup. In this setup, the training, validation, and test sets are split based on samples. All samples from all subjects are randomly shuffled and distributed into the training, validation, and test sets according to a predetermined ratio, allowing samples from the same subject to appear in three sets simultaneously. As discussed in the Preliminaries section 3, this setup has limited applicability for medical time-series-based disease diagnosis in real-world scenarios. However, it can be utilized to evaluate whether the dataset exhibits cross-subject features quickly. The results obtained from this setup are typically much higher than those from the subject-independent setup, showing a dataset’s upper limit of learnability.
Results. We evaluate the EEG dataset ADFD using this setup to provide a direct comparison of results with the subject-independent setup. The results are presented in Table 2. Our method outperforms all the baselines, achieving the top-1 results in all six evaluations, with an impressive F1 score of 97.50%. Notably, baseline methods like Informer, Nonformer, Reformer, and Transformer also demonstrate strong performance, achieving F1 scores exceeding 90%. The overall results indicate the presence of discernible and learnable features related to Alzheimer’s Disease within this dataset.
@@ 5.2 Results of Subject-Independent
Setup. In this setup, the training, validation, and test sets are split based on subjects. All subjects and their corresponding samples are distributed into the training, validation, and test sets according to a predetermined ratio or subject IDs. Samples from the same subjects should exclusively appear in one of these three sets. This setup simulates real-world medical time-series-based disease diagnosis, wherein the aim is to train a model on subjects with known labels and then test it on unseen subjects to determine if they have a specific disease. The challenges associated with this setup are discussed in section 2. All five datasets are evaluated using this setup.
@@
Results. Table 3 presents the results of the subject-independent setup. Our method achieves the top-1 F1 scores on 4 out of 5 datasets. Overall, our method achieves 15 top-1 and 30 top-3 rankings out of 30 evaluations conducted across 5 datasets and 10 baselines, considering 6 different metrics. Figure 4 provides an overview heatmap table of average rank across 5 datasets on 6 metrics for all methods. Lower rank numbers indicate better results, with rank 1 representing the best performance among all methods. Our method demonstrates the best average rank among all methods across the 6 metrics. Additionally, it is notable that the result for ADFD is a 50.65% F1 score under the subject-independent setup, which is significantly lower than the 97.50% F1 score achieved under the subject-dependent setup. This comparison highlights the challenge of the subject-independent setup, which better simulates real-world scenarios.
@@
6 Conclusion
This paper introduces Medformer, a multi-granularity patching transformer designed for medical time series classification. We develop three novel mechanisms to leverage the unique characteristics of medical time series, such as channel correlations and frequency band biomarkers. These mechanisms include cross-channel patching to learn multi-timestamp and multi-channel features, multi-granularity embedding to capture features at different scales, and two-stage multi-granularity self-attention to capture features within and among granularities. Results on five datasets, compared with ten baselines under a subject-independent setup, demonstrate the effectiveness and stability of our method, highlighting its potential for real-world applications. The limitations and future works of our method are discussed in Appendix F.2.
@@ Appendix A Data Augmentation Banks
In the embedding stage, we apply data augmentation to the patch embeddings. We utilize a bank of data augmentation techniques to enhance the model’s robustness and generalization. During the forward pass in training, each patch will pick one augmentation from available augmentation options with equal probability. The data augmentation techniques include masking, jittering, scaling, and can be further expanded to more choices. We provide a detailed description of each technique below.
A.1 Masking
Masking is a common data augmentation technique in sequence modeling tasks. It randomly masks a portion of the input sequence to encourage the model to learn from the remaining information. In our implementation, we randomly mask each element of $\mathbf{x}_{\text{emb}}^{(i)}$ with a probability of $p_{\text{mask}}$. The masked elements are replaced with zeros. The masking probability $p_{\text{mask}}$ is a hyperparameter that controls the masking rate.
A.2 Jittering
Jittering is also a common data augmentation technique by adding random noise to the input sequence. It helps the model to learn robust features by introducing variations in the input data. In our implementation, we add Gaussian noise to each element of $\mathbf{x}_{\text{emb}}^{(i)}$ with a mean of 00 and a standard deviation of $\sigma_{\text{jitter}}$. The jittering standard deviation $\sigma_{\text{jitter}}$ is a hyperparameter that controls the noise level.
A.3 Scaling
Scaling is yet another data augmentation technique that rescales the input sequence by a random factor. It helps the model to learn invariant features by adjusting the input data’s scale. In our implementation, we multiply each element of $\mathbf{x}_{\text{emb}}^{(i)}$ by a random scaling factor sampled from a Gaussian distribution with a mean of $1$ and a standard deviation of $\sigma_{\text{scale}}$. The scaling standard deviation $\sigma_{\text{scale}}$ is a hyperparameter that controls the scaling level.
@@ B.2 TDBrain Preprocessing
B.3 ADFD Preprocessing
The Alzheimer’s Disease and Frontotemporal Dementia (ADFD) dataset***https://openneuro.org/datasets/ds004504/versions/1.0.6, referenced in the papers [57, 58], is a public EEG time series dataset with 3 classes, including 36 Alzheimer’s disease (AD) patients, 23 Frontotemporal Dementia (FD) patients, and 29 healthy control (HC) subjects. The dataset has 19 channels, and the raw sampling rate is 500Hz. Each subject has a trial, with trial durations of approximately 13.5 minutes for AD subjects (min=5.1, max=21.3), 12 minutes for FD subjects (min=7.9, max=16.9), and 13.8 minutes for HC subjects (min=12.5, max=16.5). A bandpass filter between 0.5-45Hz is applied to each trial. We downsample each trial to 256Hz and segment them into non-overlapping 1-second samples with 256 timestamps, discarding any samples shorter than 1 second. This process results in 69,752 samples. For the training, validation, and test set splits, we employ both the subject-dependent and subject-independent setups. For the subject-dependent setup, we allocate 60%, 20%, and 20% of total samples into the training, validation, and test sets, respectively. For the subject-independent setup, we allocate 60%, 20%, and 20% of total subjects with their corresponding samples into the training, validation, and test sets, respectively.
@@ Appendix C Implementation Details
Medformer (Our Method) We use a list of patch lengths in patch embedding to generate patches with different granularities. Instead of flattening the patches and mapping them to dimension $D$ during patch embedding, we use a conv2d network to directly map patches into a 1-D representation with dimension $D$. These patch lengths can vary, including different numbers of patch lengths such as $\{2,4,8,16\}$, repetitive numbers such as $\{8,8,8,8\}$, or a mix of different and repetitive lengths such as $\{8,8,8,16,16,16\}$. It is also possible to use only one patch length, such as $\{8\}$, which indicates a single granularity. The patch lists used for the datasets APAVA, TDBrain, ADFD, PTB, and PTB-XL are $\{2,2,2,4,4,4,16,16,16,16,32,32,32,32,32\}$, $\{8,8,8,16,16,16\}$, $\{2,4,8,8,16,16,16,16,32,32,32,32,32,32,32,32\}$, $\{2,4,8,8,16,16,16,32,32,32,32,32\}$, and $\{2,4,8,8,16,16,16,16,32,32,32,32,32,32,32,32\}$, respectively. The data augmentations are randomly chosen from a list of four possible options: none, jitter, scale, and mask. The number following each augmentation method indicates the degree of augmentation. Detailed descriptions of these methods can be found in Appendix A. The augmentation methods used for the datasets APAVA, TDBrain, ADFD, PTB, and PTB-XL are {none, mask0.35}, {none, mask0.25}, {mask0.5}, {mask0.5}, and {jitter0.2, scale0.2, mask0.5}, respectively.
@@ D.2 Patch Length Study
Furthermore, we assess the performance of a manually selected combination of varying patch lengths, specifically $\{2,2,4,16,32\}$. This configuration achieves the highest performance across all evaluated metrics, underscoring the effectiveness of our designed attention module in accommodating multi-granularity patches. However, it is worth noting that mixing different patch lengths does not guarantee improved performance. See F for more detailed discussion.
@@
Appendix E Complexity Analysis

### v2

@@ 4 Method
Each router contains global information specific to one granularity by doing intra-granularity self-attention. By performing self-attention among routers, information can be exchanged and learned across different granularities, effectively capturing features across various scales. Additionally, the use of the router mechanism successfully reduces the time complexity of the naive approach from $O\left(\left(\sum_{i=1}^{n}N_{i}\right)^{2}\right)$ to $O\left(\sum_{i=1}^{n}N_{i}^{2}+n^{2}\right)$. Given that $N_{i}\leq T$, the worst-case time complexity for our self-attention mechanism is $O\left(nT^{2}+n^{2}\right)$. However, a reasonable choice of patch lengths as a power series, i.e., $L_{i}=2^{i}$, leads to a time complexity of $O(T^{2})$. To further reduce complexity and memory consumption, we apply shared layer normalization and feed-forward layers across all granularities. See appendix F for more details about complexity analysis.
@@ 5 Experiments
Table 2: The information of processed datasets. The table shows the number of subjects, samples, classes, channels, sampling rate, sample timestamps, modality of MedTS, and file size. Here, #-Timestamps indicates the number of timestamps per sample.
Datasets #-Subject #-Sample #-Class #-Channel #-Timestamps Sampling Rate Modality File Size
APAVA 23 5,967 2 16 256 256Hz EEG 186MB
ADFTD 88 69,752 3 19 256 256Hz EEG 2.52GB
TDBrain 72 6,240 2 33 256 256Hz EEG 571MB
PTB 198 64,356 2 15 300 250Hz ECG 2.15GB
PTB-XL 17,596 191,400 5 12 250 250Hz ECG 4.28GB
Datasets. (1) APAVA [67] is an EEG dataset where each sample is assigned a binary label indicating whether the subject has Alzheimer’s disease. (2) TDBrain [68] is an EEG dataset with a binary label assigned to each sample, indicating whether the subject has Parkinson’s disease. (3) ADFTD [69, 19] is an EEG dataset with a three-class label for each sample, categorizing the subject as Healthy, having Frontotemporal Dementia, or Alzheimer’s disease. (4) PTB [70] is an ECG dataset where each sample is labeled with a binary indicator of Myocardial Infarction. (5) PTB-XL [71] is an ECG dataset with a five-class label for each sample, representing various heart conditions. Table 2 provides information on the processed datasets. For additional details on data characteristics, train-validation-test splits under different setups, and data preprocessing, please see Appendix B.
@@
For data augmentation methods, we provide six widely used methods in time series augmentation [72, 66, 73, 61]. For more details about these six methods, see Appendix A. For the parameter tuning in our method and all baselines, we employ 6 layers for the encoder, set the dimension $D$ to 128, and the hidden dimension of feed-forward networks to 256. We utilize the Adam optimizer with a learning rate of 1e-4. The batch size is set to {32,32,128,128,128} for datasets APAVA, TDBrain, ADFTD, PTB, and PTB-XL, respectively. The training epoch is set to 100, with early stopping triggered after 10 epochs without improvement in the F1 score on the validation set. We save the model with the best F1 score on the validation set and evaluate it on the test set. See Appendix C for any additional implementation details of our method and all baselines.
Table 3: Results of Subject-Dependent Setup. The training, validation, and test sets are split based on samples according to a predetermined ratio. Results of the ADFTD dataset under this setup are presented here.
@@
ADFTD (3-Classes) Autoformer 87.83$\pm$1.62 87.63$\pm$1.66 87.22$\pm$1.97 87.38$\pm$1.79 96.59$\pm$0.88 93.82$\pm$1.64
@@
Table 4: Results of Subject-Independent Setup. The training, validation, and test sets are distributed based on subjects according to a predetermined ratio/IDs. Results of the APAVA, TDBrain, ADFTD, PTB, and PTB-XL datasets under this setup are presented here.
@@
ADFTD (3-Classes) Autoformer 45.25$\pm$1.48 43.67$\pm$1.94 42.96$\pm$2.03 42.59$\pm$1.85 61.02$\pm$1.82 43.10$\pm$2.30
@@ 5.1 Results of Subject-Dependent
Setup. In this setup, the training, validation, and test sets are split based on samples. All samples from all subjects are randomly shuffled and distributed into the training, validation, and test sets according to a predetermined ratio, allowing samples from the same subject to appear in three sets simultaneously. As discussed in the Preliminaries section 3, this setup has limited applicability for MedTS-based disease diagnosis in real-world scenarios. It is usually used to evaluate whether the dataset exhibits cross-subject features quickly. The results obtained from this setup are typically much higher than those from the subject-independent setup, showing a dataset’s upper limit of learnability.
Results. We evaluate the EEG dataset ADFTD using this setup to provide a direct comparison of results with the subject-independent setup. The results are presented in Table 3. Our method outperforms all the baselines, achieving the top-1 results in all six evaluations, with an impressive F1 score of 97.50%. Notably, baseline methods like Informer, Nonformer, Reformer, and Transformer also demonstrate strong performance, achieving F1 scores exceeding 90%. The overall results indicate the presence of discernible and learnable features related to Alzheimer’s Disease within this dataset.
@@ 5.2 Results of Subject-Independent
Setup. In this setup, the training, validation, and test sets are split based on subjects. All subjects and their corresponding samples are distributed into the training, validation, and test sets according to a predetermined ratio or subject IDs. Samples from the same subjects should exclusively appear in one of these three sets. This setup simulates real-world MedTS-based disease diagnosis, aiming to train a model on subjects with known labels and then test it on unseen subjects to determine if they have a specific disease. The challenges associated with this setup are discussed in section 2. All five datasets are evaluated using this setup.
@@
Results. Table 4 presents the results of the subject-independent setup. Our method achieves the top-1 F1 scores on 4 out of 5 datasets. Overall, our method achieves 15 top-1 and 30 top-3 rankings out of 30 evaluations conducted across 5 datasets and 10 baselines, considering 6 different metrics. Figure 4 provides an overview heatmap table of average rank across 5 datasets on 6 metrics for all methods. Lower rank numbers indicate better results, with rank 1 representing the best performance among all methods. Our method demonstrates the best average rank among all methods across the 6 metrics. Additionally, it is notable that the result for ADFTD is a 50.65% F1 score under the subject-independent setup, which is significantly lower than the 97.50% F1 score achieved under the subject-dependent setup. This comparison highlights the challenge of the subject-independent setup, which better simulates real-world scenarios.
5.3 Ablation Study and Additional Experiments
@@
Additional Experiments. We also perform experiments on two human activities recognition datasets [74, 75] to demonstrate the learning ability of our model on general time series with potential channel correlations(Appendix E).
6 Conclusion and Limitations
Conclusion This paper presents Medformer, a multi-granularity patching transformer tailored for MedTS classification. We introduce three novel mechanisms that leverage the distinctive features of MedTS, including channel correlations and frequency band biomarkers. These mechanisms include cross-channel patching to capture multi-timestamp and cross-channel features, multi-granularity embedding to learn features at various scales, and a two-stage multi-granularity self-attention mechanism to extract features both within and across granularities. Experimental results across five datasets, evaluated against ten baselines under the subject-independent setup, demonstrate the effectiveness and robustness of our approach, underscoring its potential for real-world applications.
Limitations and Future Work The design of Medformer allows for inputting various patch lengths, offering both opportunities and challenges. While variable patch lengths provide flexibility and have been shown to outperform uniform lengths in some cases (see Appendix D.2 and Appendix C), not all patch length configurations yield optimal results. Some combinations may perform worse than uniform patch lengths, necessitating careful tuning of patch lengths as hyperparameters. Future work could explore mechanisms for automatically selecting the most effective patch lengths and optimizing the model’s ability to capture relevant granularities. Additionally, while our method outperforms baselines and performs well on certain datasets under the subject-independent setup, it does not include mechanisms specifically designed for this setup. Developing modules that decompose subject-specific features from task-related features could further enhance learning under the subject-independent setup, presenting an intriguing direction for future research.
@@ Appendix A Data Augmentation Banks
In the embedding stage, we apply data augmentation to the patch embeddings. We utilize a bank of data augmentation techniques to enhance the model’s robustness and generalization. During the forward pass in training, each patch will pick one augmentation from available augmentation options with equal probability. The data augmentation methods include temporal flipping, channel shuffling, temporal masking, frequency masking, jittering, and dropout, and can be further expanded to more choices. We provide a detailed description of each technique below.
Temporal Flippling We reverse the MedTS data along the temporal dimension. The probability of applying this augmentation is controlled by a parameter prob, with a default value of 0.5.
Channel Shuffling We randomly shuffle the order of MedTS channels. The probability of applying channel shuffling is controlled by the parameter prob, also set by default to 0.5.
temporal masking We randomly mask some timestamps across all channels. The proportion of timestamps masked is controlled by the parameter ratio, with a default value of 0.1.
Frequency Masking First introduced in [73] for contrastive learning, this method involves converting the MedTS data into the frequency domain, randomly masking some frequency bands, and then converting it back. The proportion of frequency bands masked is controlled by the parameter ratio, with a default value of 0.1.
Jittering Random noise, ranging from 0 to 1, is added to the raw data. The intensity of the noise is adjusted by the parameter scale, which is set by default to 0.1.
Dropout Similar to the dropout layer in neural networks, this method randomly drops some values. The proportion of values dropped is controlled by the parameter ratio, with a default setting of 0.1.
@@ B.2 TDBrain Preprocessing
B.3 ADFTD Preprocessing
The Alzheimer’s Disease and FronTotemporal Dementia (ADFTD) dataset***https://openneuro.org/datasets/ds004504/versions/1.0.6, referenced in the papers [69, 19], is a public EEG time series dataset with 3 classes, including 36 Alzheimer’s disease (AD) patients, 23 Frontotemporal Dementia (FTD) patients, and 29 healthy control (HC) subjects. The dataset has 19 channels, and the raw sampling rate is 500Hz. Each subject has a trial, with trial durations of approximately 13.5 minutes for AD subjects (min=5.1, max=21.3), 12 minutes for FD subjects (min=7.9, max=16.9), and 13.8 minutes for HC subjects (min=12.5, max=16.5). A bandpass filter between 0.5-45Hz is applied to each trial. We downsample each trial to 256Hz and segment them into non-overlapping 1-second samples with 256 timestamps, discarding any samples shorter than 1 second. This process results in 69,752 samples. For the training, validation, and test set splits, we employ both the subject-dependent and subject-independent setups. For the subject-dependent setup, we allocate 60%, 20%, and 20% of total samples into the training, validation, and test sets, respectively. For the subject-independent setup, we allocate 60%, 20%, and 20% of total subjects with their corresponding samples into the training, validation, and test sets, respectively.
@@ Appendix C Implementation Details
Medformer (Our Method) We use a list of patch lengths in patch embedding to generate patches with different granularities. Instead of flattening the patches and mapping them to dimension $D$ during patch embedding, we use a conv2d network to directly map patches into a 1-D representation with dimension $D$. These patch lengths can vary, including different numbers of patch lengths such as $\{2,4,8,16\}$, repetitive numbers such as $\{8,8,8,8\}$, or a mix of different and repetitive lengths such as $\{8,8,8,16,16,16\}$. It is also possible to use only one patch length, such as $\{8\}$, which indicates a single granularity. The patch lists used for the datasets APAVA, TDBrain, ADFD, PTB, and PTB-XL are $\{2,2,2,4,4,4,16,16,16,16,32,32,32,32,32\}$, $\{8,8,8,16,16,16\}$, $\{2,4,8,8,16,16,16,16,32,32,32,32,32,32,32,32\}$, $\{2,4,8,8,16,16,16,32,32,32,32,32\}$, and $\{2,4,8,8,16,16,16,16,32,32,32,32,32,32,32,32\}$, respectively. The data augmentations are randomly chosen from a list of four possible options: none, jitter, scale, and mask. The number following each augmentation method indicates the degree of augmentation. Detailed descriptions of these methods can be found in Appendix A. The augmentation methods used for the datasets APAVA, TDBrain, ADFD, PTB, and PTB-XL are {none, drop0.35}, {none, drop0.25}, {drop0.5}, {drop0.5}, and {jitter0.2, scale0.2, drop0.5}, respectively.
@@ D.2 Patch Length Study
Furthermore, we assess the performance of a manually selected combination of varying patch lengths, specifically $\{2,2,4,16,32\}$. This configuration achieves the highest performance across all evaluated metrics, underscoring the effectiveness of our designed attention module in accommodating multi-granularity patches. However, it is worth noting that mixing different patch lengths does not guarantee improved performance. See G for more detailed discussion.
@@
Appendix E Additional Experiments
Table 7: Additional Datasets and Methods We selected three old baselines in the previous experiments that showed strong performance: Crossformer, Reformer, and Transformer. Additionally, we introduce three new baselines: TCN, ModernTCN, and Mamba. These six baselines are evaluated on one old dataset in the previous experiments, TDBrain(6,240 samples, 2 classes), and two new HAR datasets: FLAAP (13,123 samples, 10 classes) and UCI-HAR (10,299 samples, 6 classes). The bold number denotes the best result, and the underlined number denotes the second best.
Datasets TDBrain (2 Classes) FLAAP (10 Classes) UCI-HAR (6 Classes)
Models Metrics Accuracy F1 Score Accuracy F1 Score Accuracy F1 Score
[…]
Reformer 87.92$\pm$2.01 87.85$\pm$2.08 71.65$\pm$1.27 71.14$\pm$1.45 88.44$\pm$2.02 88.34$\pm$1.98
Transformer 87.17$\pm$1.67 87.10$\pm$1.68 74.96$\pm$1.25 74.49$\pm$1.39 88.86$\pm$1.65 88.80$\pm$1.67
TCN 80.92$\pm$2.94 80.82$\pm$3.03 66.48$\pm$1.66 65.29$\pm$1.74 93.08$\pm$0.95 93.19$\pm$0.88
ModernTCN 81.96$\pm$2.12 81.79$\pm$2.23 74.80$\pm$0.96 74.35$\pm$0.85 91.44$\pm$1.01 91.47$\pm$0.98
[…]
To evaluate the performance of our method on general time series, we test it on two human activity recognition (HAR) datasets: FLAAP [74] and UCI-HAR [75], which exhibit potential channel correlations inherently. Additionally, we compare our method with three other approaches: TCN [51], ModernTCN [16], and Mamba [77]. Our method achieves the highest top-1 accuracy and F1 score on TDBrain and FLAAP, and ranks second-best on UCI-HAR.
Appendix F Complexity Analysis